what does b equal if the equation is
69=61/108(105/2)+b

Answer by checkley71(8403)

69=(61/108)(105/2)+B
69=(61*105)/(108*2)+B
69=(6405/216)+B
69=29.653+B
B=69-29.653
B=39.347 ANSWER.